OpenStudy (anonymous):

Form a polynomial F(x) with real coefficients having the given degree and zeros Degree: 4; zeros -5+3i; 4 multiplicity 2

OpenStudy (anonymous):

ick

OpenStudy (anonymous):

4 with multiplicity 2 means one factor is \((x-4)^2\) then you have to find a quadratic with zeros \(-5+3i\) do you know how to do that? it is not that hard

OpenStudy (anonymous):

turn the roots into factor form of the polynomial

OpenStudy (anonymous):

unfortunately Im not sure how to find the zeros, I've checked examples, but I havnt grasped how though :(

OpenStudy (anonymous):

hold the phone

OpenStudy (anonymous):

you are not being asked to find the zeros, you are TOLD the zeros you have to come up with the polynomial

OpenStudy (anonymous):

you have "zeros -5+3i; 4 multiplicity 2 "

OpenStudy (anonymous):

sorry I meant find a quadractic with zeros

OpenStudy (anonymous):

"4 multiplicity 2" means it has two factors of \(x-4\) i.e. it has a factor of \((x-4)^2\) here is a simple method to find the quadratic work backwards, it requires a little algebra, then i can show you a method that requires only memorizaton

OpenStudy (anonymous):

put \[x=-5+3i\] work backwards by adding \(5\) to get \[x+5=3i\] then square both sides (carefully) and get \[x^2+10x+25=(3i)^2=-9\] then add \(9\) and get \[x^2+10x+34=0\]

OpenStudy (anonymous):

so the quadratic with zeros \(-5+3i\) and \(-5-3i\) is \(x^2+10x+34\)

OpenStudy (anonymous):

if you want a really really snap way, that requires no algebra but it does require memorizing, is that the quadratic with zeros \(a+bi\) is \[x^2-2ax+(a^2+b^2)\]

OpenStudy (anonymous):

in your case you had \(-5+3i\) so \(a=-5,b=3\) and the quadratic is \[x^2-2\times(-5)x+(5^2+9^2)=x^2+10x+34\]

OpenStudy (anonymous):

ok I have a question, when you said square both sides (carefully), where did you get the 10x from?

OpenStudy (anonymous):

and your final job is to multiply out \[(x^2+10x+34)(x-4)^2\] i would cheat so as not to screw up the algebra

OpenStudy (anonymous):

lol that is exactly what i meant by "carefully"

OpenStudy (anonymous):

\((x+5)^2=(x+5)(x+5)=x^2+10x+25\) and not for example \(x^2+25\)

OpenStudy (anonymous):

ohh lol ok I understand now

OpenStudy (anonymous):

Thank you so much satelite73!

OpenStudy (anonymous):

yw you still have to multiply all that muck out "carefully"!

OpenStudy (anonymous):

haha oh I will!

OpenStudy (anonymous):

if you get this http://www.wolframalpha.com/input/?i=%28x^2%2B10x%2B34%29%28x-4%29^2 you know you are right

OpenStudy (anonymous):

Ok cool, i'll check it out then